An experiment requires a solution which is 98 mM glucose and 7 mM MgSO4. You have distilled water, a 3.4 M stock solution of glucose, and a 0.43 M MgSO4 stock solution in a lab. The experiment requires 10 ml of the solution in a test tube.

(a) What volume of the stock glucose solution do you need in the test tube?

(b) What volume of stock MgSO4 solution do you need in the test tube?
1
Expert's answer
2018-11-20T15:44:41-0500

Answer on Question #83068, Chemistry/ General Chemistry

An experiment requires a solution which is 98 mM glucose and 7 mM MgSO4. You have distilled water, a 3.4 M stock solution of glucose, and a 0.43 M MgSO4 stock solution in a lab. The experiment requires 10 ml of the solution in a test tube.


(a) What volume of the stock glucose solution do you need in the test tube?


(b) What volume of stock MgSO4 solution do you need in the test tube?

Solution

a) Find number of moles of glucose in 10 mL of the solution in a test tube, the concentration of glucose in this solution is 9810-3 mol/L

n(glucose) = cV = 9810-3 mol/L  0.01 L = 9.810-4 mol

Find volume of 3.4M stock solution of glucose:

V=n/c

V= 9.810-4 mol / 3.4 mol/L = 2.8810-4 L = 0.288 mL

b) Find number of moles of MgSO4 in 10 mL of the solution in a test tube, the concentration of MgSO4 in this solution is 710-3 mol/L

n(MgSO4) = cV = 710-3 mol/L  0.01 L = 710-5 mol

Find volume of 0.43 M stock solution of MgSO4:

V=n/c

V= 710-5 mol / 0.43 mol/L = 1.6310-4 L = 0.163 mL

To prepare 10 mL of solution which is 98 mM glucose and 7 mM MgSO4 one needs to take 0.288 mL of 3.4 M stock solution of glucose and 0.163 mL of 0.43 M stock solution of MgSO4, then one should add necessary volume of distilled water to reach 10 mL line.

Answer: a) 0.288 mL of 3.4 M stock solution of glucose

b)0.163 mL of 0.43M stock solution of MgSO4